Calculate amount (in mol) of magnesium and oxygen
1
Expert's answer
2020-05-07T14:08:46-0400

Magnesium reacts with oxygen to produce magnesium oxide:

2Mg(s) + O2(g) ----> 2MgO(s)

Number of moles of Mg = mass(Mg) / Mr(Mg) = mass(Mg) / 24.3

Number of moles of O2 = mass(O2) / Mr(O2) = mass(O2) / 32.0